Find equation: (2,5) (-6,-5)
I know that you use m=y2-y1/ x2-x1
which is 5-(-5)= 10 5/4 then: y=5/4x +b
2-(-6)= 8 5= 5/4 +b
5= 5/4(2)+b
5= 5/2 +b
5-5/2=b here's where it gets confusing
the answer is.... the book states y= 5/4x+5/2

Answer by checkley77(12844)

(2,5)(-6,-5)
m=(-5-5)/-6-2)
m=-10/-8=5/4 is the slope.
Y=mX+b
5=5/4*2+b
5=10/4+b
b=5-10/4
b=5-2.5
b=2.5 thge y intercept.
y=5x/4-2.5 answer. Same answer in the book.